Find the equation of the line through points (-3, 6) and (10, 12).
Answer by Alan3354(69443)

Find the equation of the line through points (-3, 6) and (10, 12).
---------
Use determinants:
|+x +y +1|
|-3 +6 +1| = 0
|10 12 +1|
---------
x(6-12) -y(-3-10) +(-36-60) = 0
-6x + 13y = 96
6x - 13y = -96